If the Kb of a weak base is 4.9 × 10-6, what is the pH of a 0.49 M solution of this base?
Answer #1

Consider a weak base MOH.

MOH---->M++OH-

Kb=[M+][OH-]​​​​​​​/[MOH]

We know that [M+]=[OH-]

Kb= [OH]2/[MOH]

[MOH]=0.49(given)

Kb=4.9×10-6

[OH]2=Kb×[MOH] =4.9×10-6×0.49=2.4×10-6.

[OH]=√2.4×10-6M =1.5×10-3

pOH= -log([OH-])=-log(1.5×10-3) =2.80

Therefore pH = 14-2.80=11. 2

pH=11.2

As pH + pOH =14.